# Bridging the Gap Between LLMs and Human Intentions: Progresses and Challenges in Instruction Understanding, Intention Reasoning, and Reliable Generation

### v1

@@ Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation. However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of reliable generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
@@ 1 Introduction
Figure 1: Example of large language model generation.
I. Challenge of Instruction Understanding. One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al. (2024), and context-aware optimization Sun et al. (2024). While these methods show promise, they often fall short when addressing the complexities and ambiguities present in real-world instructions.
II. Challenge of Intention Reasoning. Another critical area is intention reasoning as illustrated in Figure 1(b), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
III. Challenge of Reliable Generation. The final major challenge is the reliable generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c). While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issue involve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
@@
Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020)
2) Attention OptimizationChen et al. (2024)
@@
Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ 1) SFT Teng et al. (2024); Sun et al. (2024), 2) RL Zhou et al. (2024) , understanding_leaf, text width=32em ] ] ] [ Intention
@@
Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2023),
CAR Weller et al. (2024) , reasoing_leaf, text width=32em ] ] [ Fuzzy Language
Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Folkscope Yu et al. (2023), Miko Lu et al. (2024), ATC Deng et al. (2023) , reasoing_leaf, text width=32em ] ] [ Intention
@@
Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=32em ] ] ] [ Reliable
@@
ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al. (2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al. (2024), CalibrateMath Lin et al. (2022) , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
@@
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable content generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al. (2024); Plaat et al. (2024); Huang et al. (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, and generate content that aligns with human intentions, minimizing hallucinations and ensuring credibility and ethical consistency.
@@
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. Lou et al. (2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Gao et al. analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions. Xu et al. Xu et al. (2024b) examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. Plaat et al. (2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. Shorinwa et al. (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, with a unique classification and systematic analysis of instruction processing, addressing current solutions to key challenges.
@@
As in Figure 2, we begin by exploring the capability of user instruction understanding (§2). Next, we analyze intention comprehension and reasoning, focusing on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3). We then examine reliable content generation, assessing the quality and credibility of model-generated outputs (§4). Following this, we review benchmarks that address these core challenges (§5). Finally, we propose potential research directions (§6) and summarize the key findings (§7).
Figure 3: Common understanding issues in LLMs include: (A) Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context (A.1). (B) Incorrect Relevance Judgment(§2.2), such as the model incorrectly associates wrong content (B.1-2.nd) from the previous turn (B.1-1.st).
@@ 2.1 Long-Text Comprehension
Challenges remain in understanding long context and this paper classifies the factors into three categories as follows:1) Information Sparsity and Redundancy. Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction and potential hallucinations. 2) Remote Information Failure (Figure 3 A.1). Long contexts may cause models to forget relevant information that is distant within the text. Additionally, links between remote information across paragraphs or sentences can be difficult for models to identify, diminishing their understanding of contextual connections. 3) Attention Dilution. As context length increases, the model’s attention mechanism faces greater computational demands and struggles to assign appropriate weights to each token, making it harder to prioritize key information, particularly with complex, multi-level relationships in longer texts. This paper classifies the existing solutions into the following two categories:
@@ 2.2 Multi-Turn Conversation Handling
This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al. (2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al. (2024). Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al. (2024). 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al. (2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al. (2024). 3) Incorrect Relevance Judgment (Figure 3 B.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al. (2024).
@@ 3 Intention Reasoning
Figure 4: Common reasoning issues in LLMs include: (A) Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (A.1) or overlooking logical inconsistencies (A.2), (B) Misinformation Reasoning (§3.2), caused by temporal misalignment leading to outdated responses (B.1) or data contamination resulting in misleading outputs (B.2). (C) Fuzzy Language Interpretation (§3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2). and (D) Intention Clarification Failure (§3.4) , where it misinterprets sarcasm (D.1) or ignores prior emotional context (D.2).
@@ 3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks Li et al. (2024a); Zheng et al. (2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 4 A.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 4 A.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
@@
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 Jin et al. (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods. MacNoise Hong et al. (2024) enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination.
@@ 3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al. (2023); Xu et al. (2024a). From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 4 B.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 4 B.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL Jang et al. (2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al. (2022). BIPIA Yi et al. (2023) used adversarial training to combat the effects of information pollution and improve model robustness. CAR Weller et al. (2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.

### v2

# A Survey of Progress and Challenges in Aligning LLMs with Human Intentions

@@ Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation. However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and Reliable Dialog Generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of Reliable Dialog Generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
A Survey of Progress and Challenges in Aligning LLMs
with Human Intentions
Zongyu Chang1thanks: These authors contributed equally to this work., Feihong Lu2*, Ziqin Zhu2, Qian Li1, Cheng Ji2, Zhuo Chen1, Hao Peng2, Yang Liu5, Ruifeng Xu3, Yangqiu Song4, Jianxin Li2, Shangguang Wang1 1Beijing University of Posts and Telecommunications, 2Beihang University, 3Harbin Institute of Technology, Shenzhen, 4Hong Kong University of Science and Technology, 5Chinese Academy of Sciences {changzongyu, li.qian, sgwang}@bupt.edu.cn, {lufeihong, zhuziqin, jicheng, lijx}@buaa.edu.cn, xuruifeng@hit.edu.cn, yqsong@cse.ust.hk
@@ 1 Introduction
Figure 1: Example of LLMs generation.
Figure 2: Unlike previous surveys on LLMs, we do not consider the alignment between LLMs and humans as an isolated process,instead, we view it as a continuous and dynamic information processing process consisting of instruction understanding, intention reasoning, and reliable dialogue generation.
I. Challenge of Instruction Understanding. One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a) and Figure (2I), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al. (2024), and context-aware optimization Sun et al. (2024). While these methods show promise, they often fall short when addressing the complexities and ambiguities present in instructions.
II. Challenge of Intention Reasoning. Another critical area is intention reasoning as illustrated in Figure 1(b) and Figure 2 (II), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
III. Challenge of Reliable Dialog Generation. The final major challenge is the reliable dialog generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c) and Figure 2 (III. While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issue involve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
@@
Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020),
2) Attention Optimization Chen et al. (2024),
@@
Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Soft
Fine-tuning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ Multi-turn SFT Teng et al. (2024), Parrot Sun et al. (2024) , understanding_leaf, text width=26.5em ] ] [ Reforcement
Learning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ ArCHer Zhou et al. (2024) , understanding_leaf, text width=26.5em ] ] ] ] [ Intention
@@
Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Targeted
Fine-tuning, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2025),
KC-LLMs Wang et al. (2023), CAR Weller et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Fuzzy Language
Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Clue
Engineering , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ Folkscope Yu et al. (2023), Miko Lu et al. (2024), ATC Deng et al. (2023) , reasoing_leaf, text width=26.5em ] ] ] [ Intention
@@
Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Deep
Reasoning , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] ] [ Reliable
@@
ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al. (2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al. (2024), CalibrateMath Lin et al. , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
@@
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable dialog generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al. (2024); Plaat et al. (2024); Huang et al. (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, infer user intentions, and generate content that closely with human intentions, thereby maximizing the alignment between LLMs and humans.
@@
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. Lou et al. (2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Gao et al. analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions. Xu et al. Xu et al. examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. Plaat et al. (2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. Shorinwa et al. (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, as shown in Figure 2 and Figure 3, with a unique classification and systematic analysis of instruction processing, while addressing current solutions to key challenges.
@@
As in Figure 3, we begin by exploring the capability of user instruction understanding (§2). Next, we focu on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3). We then examine reliable dialog generation, assessing the quality and credibility of model-generated outputs (§4). Next, we briefly analyze the problems faced by LLMs in face of different challenges(§5) and review the benchmarks(§6) for the above problems. Finally, we propose potential research directions (§7) and summarize the key findings (§8).
@@ 2.1 Long-Text Comprehension
Understanding lengthy textual instructions remains a significant hurdle for large language models (LLMs), as real-world human instructions are often expressed in loose, unstructured natural language, contrasting with the explicitly defined tasks and structured labeling commonly employed in cue word engineering, so we categorize the relevant factors into the following three categories:1) Information Sparsity and Redundancy. Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction. 2) Remote Information Failure (Figure 4). Long contexts may cause models to forget relevant information that is distant within the text. Additionally, links between remote information across paragraphs or sentences can be difficult for models to identify, diminishing their understanding of contextual connections. 3) Attention Dilution. As context length increases, the model’s attention mechanism faces greater computational demands and struggles to assign appropriate weights to each token, making it harder to prioritize key information, particularly with complex, multi-level relationships in longer texts. This paper classifies the existing solutions into the following two categories:
@@ 2.2 Multi-Turn Conversation Handling
Multi-turn conversation serves as a fundamental interaction mode between LLMs and humans. Given the challenges users face in providing complete and precise instructions in a single turn, they often opt to refine and clarify their intentions incrementally through iterative exchanges. However, due to the characteristics of real-world conversations, such as the constant changes in user intentions and long-distance dependencies, LLMs still faces significant challenges in achieving coordinated multi-round interactions with humans.This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al. (2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al. (2024). Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al. (2024). 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al. (2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al. (2024). 3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al. (2024).
@@ 3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks Li et al. (2024a); Zheng et al. (2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 6 Q.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
@@
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 Jin et al. (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods.
@@ 3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al. (2023); Xu et al. (2024). From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 7 Q.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 7 Q.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL Jang et al. (2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al. (2022). RKC-LLMs Wang et al. (2023) allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information. BIPIA Yi et al. (2025) used adversarial training to combat the effects of information pollution and improve model robustness. CAR Weller et al. (2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.